Jarstrom, Inc., is considering a project that will require an initial investment of $10,000 and will...

  1. Jarstrom, Inc., is considering a project that will require an initial investment of $10,000 and will return a perpetual cash flow of $1,200. Compute its NPV assuming k = .10.

    a. 0
    b. $2,000 c. $8,800 d. $10,000

  2. Suppose you are considering a project with $10,000 initial investment. You estimate that the project will generate a level cash flow of $1,200 per year for 15 years. At the end of the 15th year, the project will be liquidated for $3,000. Find the NPV if cost of capital is 12 percent.

    a. -$1,278.87 b. -$3,278.87 c. -$5,278.87 d. -$6,820.00

  3. Compute the yield to maturity for the FGH Company bond maturing in 20 years (7 percent coupon rate and a face value of $1,000) with a market price of 1000.

    a. 7.00% b. 8.74% c. 10.27% d. 12.93%

  4. A project costs $200 today. If a cash flow of $250 will be received at the end of seven years, what is its IRR?

    a. 1.85% b. 2.49% c. 3.24% d. 25%
Answer #1

3) NPV = Perpetual Cash flow/ Cost of equity - Investment = 1200/0.1 - 10,000 = 2000

4) NPV = 1200*(1-(1+r)-n)/r + 3000/(1+r)n - 10000  = 1200*(1-(1+12%)-15/12% + 3000/(1+12%)15 -10000 = - 1278.87

5)Since bond is sold at par value so YTM and Coupon rate are same. YTM =7%

6) 200 = 250/(1+IRR)7
IRR = (250/200)1/7-1 = 3.24%

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well